OpenStudy (anonymous):

k=1 sigma k tends to infinity (-3/2)^(k+1) the seires is converge or diverge how can i indentify this series is converge or diverge.

OpenStudy (turingtest):

ratio test?

OpenStudy (anonymous):

no ..what is ratio test:(

OpenStudy (turingtest):

if the series is \(\{a_n\}\) then the series converges absolutely (and hence also converges) if\[\lim_{n\to\infty}\left|{\{a_{n+1}\}\over\{a_n\}}\right|<1\]

OpenStudy (anonymous):

this seris is not geometric series..?

OpenStudy (turingtest):

I suppose, I just thought of one way to test for convergence but we can do it the geometric way...

OpenStudy (turingtest):

we can reewrite this\[\sum_{k=1}^\infty(-\frac32)^{k+1}=\sum_{k=1}^\infty(-\frac32)^2(-\frac32)^{k-1}\]in general, geometric series of the form\[\sum_{k=1}^\infty ar^{k-1}\]converge when \(|r|<1\) we get our answer

OpenStudy (turingtest):

we could easily find what it actually converges to as well, but that seems not to be required

OpenStudy (anonymous):

how can i find this seris sums

OpenStudy (anonymous):

heloow turing test r u ther..?

OpenStudy (anonymous):

ny one can see my question..?